Should print UDKAST on the first three pages, light.

\usepackage[light,firsttwo]{draftcopy}
\draftcopyLastPage{3}

Test4: This is a test page 1

This is a testpage 1 Test4: This is a test page 2 This is a testpage 2 Test4: This is a test page 3 This is a testpage 3 Test4: This is a test page 4 This is a testpage 4